# DISCO: Distribution-Aware Calibration for Object Detection with Noisy Bounding Boxes

### v2

@@ 5 Conclusions
In this paper, we focus on an under-explored and challenging problem termed object detection with noisy bounding boxes. Motivated by the observation about proposal aggregation, we propose DISCO to calibrate the corrupt supervision signals. Spatial distribution modeling is performed and then three distribution-aware techniques (i.e., DA-Aug, DA-Ref, and DA-Est) are adopted successively. Experiments show that our DISCO can achieve state-of-the-art performance. We believe that DISCO can serve as a stronger baseline for this task and expect it can motivate more future works in the field of object detection and learning with noisy labels.
Appendix
@@ Appendix A Details of the Experimental Setup
A.1 Noise Simulation
Following Liu et al. [2022], clean annotations are perturbed to simulate noisy bounding box annotations in our experiments, which is performed once for each dataset. Specifically, let $c_{\text{x}},c_{\text{y}},w,h$ represent the central x-axis coordinate, central y-axis coordinate, width, and height of a clean bounding box, respectively. We simulate a noisy bounding box by randomly shifting and scaling a clean one, which can be formulated as
@@ A.1 Noise Simulation
where $\Delta_{\text{x}}$, $\Delta_{\text{y}}$, $\Delta_{\text{w}}$, and $\Delta_{\text{h}}$ obey the uniform distribution $U(-n,n)$ and $n$ is the noise level. For example, when $n$ is set to $40\%$, $\Delta_{\text{x}}$, $\Delta_{\text{y}}$, $\Delta_{\text{w}}$, and $\Delta_{\text{h}}$ would ranges from $-0.4$ to $0.4$. Note that Equation 20 is conducted on each bounding box of the training set. Such a noise simulation can guarantee access to real ground-truths for analyzing training behaviors and evaluating the performance of box refinement.
A.2 Hyperparameter Selections
@@
Appendix B More Ablation Studies
In this section, we conduct more ablation studies to further verify the effectiveness of the proposed DISCO. These ablation studies contain backbone compatibility, sensitivity analysis of other hyperparameters, and the execution number of DISCO. Unless otherwise specified, the following experiments are all based on VOC at the $40\%$ noise level.
B.1 Backbone Compatibility
As mentioned in Section 4.1, following Liu et al. [2022], the benchmark experiments are performed with ResNet-50 He et al. [2016] as the backbone. To further demonstrate the superior performance of our method, we conduct an additional experiment based on different backbones. Specifically, in this experiment, DISCO is compared to OA-MIL on COCO at the $40\%$ noise level with the backbone set to ResNet-101 He et al. [2016] and Swin-T Liu et al. [2021], and other experiment setups remain the same. In this way, we aim to evaluate the performance of our DISCO for a large-scale dataset when it is equipped with an advanced backbone. The experimental results are reported in Table 7. For ResNet-101, it can be observed that our DISCO can further improve performance and still achieve state-of-the-art results. For Swin-T, although the noisy setting and training configurations may not be suitable for this backbone, compared to OA-MIL, our DISCO can still attain superior detection performance.
@@ B.1 Backbone Compatibility
Table 7: Ablation studies of backbone compatibility. The experiment is conducted on COCO at 40% noise level with ResNet-101 and Swin-T. DISCO can still outperform OA-MIL when equipped with different backbones.
B.2 Other Hyperparameter Sensitivity
We perform sensitivity analysis for other hyperparameters, including $N^{\prime}$ of Equation 9 and $\gamma,\lambda$ of Equation 19. Note that we choose some moderate values rather than extreme ones to reasonably evaluate the sensitivity for each hyperparameter. The experimental results are reported in Table 8. It can be observed that the augmented proposal number $N^{\prime}$ is insensitive when varying from $5$ to $20$. This is the reason why we empirically fix $N^{\prime}$ to $10$ for all settings. Besides, two loss weights $\gamma,\lambda$ also remain insensitive while $\lambda$ is relatively crucial. This is because it controls the strength of an extra classification loss term, directly affecting classification accuracy.
B.3 Execution Number of DISCO
Hyper. Value $\text{AP}_{50}$bigstrut
@@ B.3 Execution Number of DISCO
Hyper. Value $\text{AP}_{50}$bigstrut
[…]
0.3 68.7
[…]
Execution Number of DISCO $\text{AP}_{50}$bigstrut
[…]
2 68.7
[…]
Table 9: Ablation studies of the execution number of DISCO. Our execution strategy can achieve superior performance.

### v3

@@ 5 Conclusions
In this paper, we focus on an under-explored and challenging problem termed object detection with noisy bounding boxes. Motivated by the observation about proposal aggregation, we propose DISCO to calibrate the corrupt supervision signals. Spatial distribution modeling is performed and then three distribution-aware techniques (i.e., DA-Aug, DA-Ref, and DA-Est) are adopted successively. Experiments show that DISCO can achieve SOTA performance. We believe that DISCO can serve as a stronger baseline for this task and expect it can motivate more works in the fields of object detection and noisy label learning.
Acknowledgements
The work described in this paper was supported in part by the grants from China National Key R&D Program (Grant No. 2023YFE0202700, 2022YFE0200700), Hong Kong Innovation and Technology Fund (Project No. MHP/092/22), NSFC General Project (62072452), Regional Joint Fund of Guangdong (Grant No. 2021B1515120011), National Natural Science Foundation of China (Project No. 6237073934, 3234101132), and Natural Science Foundation of Guangdong Province (2022A1515011579).
@@ Appendix A Details of the Experimental Setup
A.1 Datasets
Two large-scale image datasets are adopted in our experiments, including Pascal VOC 2007 [Everingham et al.(2010)Everingham, Van Gool, Williams, Winn, and Zisserman] and MS-COCO 2017 [Lin et al.(2014)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]. Pascal VOC 2007 (VOC) is a standard dataset for object detection, consisting of $9,963$ images with $24,640$ box annotations. MS-COCO 2017 (COCO) is also a popular object detection benchmark, containing $328,000$ images of generic objects. Following [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang], clean annotations are perturbed to simulate noisy bounding box annotations in our experiments, which is performed once for each dataset. Specifically, let $c_{\text{x}},c_{\text{y}},w,h$ represent the central x-axis coordinate, central y-axis coordinate, width, and height of a clean bounding box, respectively. We simulate a noisy bounding box by randomly shifting and scaling a clean one, which can be formulated as
@@ A.1 Datasets
where $\Delta_{\text{x}}$, $\Delta_{\text{y}}$, $\Delta_{\text{w}}$, and $\Delta_{\text{h}}$ obey the uniform distribution $U(-n,n)$ and $n$ is the noise level. For example, when $n$ is set to $40\%$, $\Delta_{\text{x}}$, $\Delta_{\text{y}}$, $\Delta_{\text{w}}$, and $\Delta_{\text{h}}$ would ranges from $-0.4$ to $0.4$. Note that Equation 15 is conducted on each bounding box of the training set. Such a noise simulation can guarantee access to real ground-truths for analyzing training behaviors and evaluating the performance of box refinement. Noise levels are set to $\{10\%,20\%,30\%,40\%\}$ for VOC and $\{20\%,40\%\}$ for COCO.
A.2 Implementation Details
Following [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang], we implement our method on FasterRCNN [Ren et al.(2015)Ren, He, Girshick, and Sun] with ResNet-50 [He et al.(2016)He, Zhang, Ren, and Sun] as the backbone. The idea of DISCO can be easily generalized to other frameworks and we choose to perform our experiments with FasterRCNN as it is widely adopted [Kaur and Singh(2023)]. As a common practice, the model is trained with the “1×1$\times$1 ×" schedule [Girshick et al.(2018)Girshick, Radosavovic, Gkioxari, Dollár, and He]. Notably, all other training configurations are aligned with [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang] to ensure fairness. As commonly done, mean average precision (mAP@.5.5.5.5) and mAP@[.$[.5,95]$ are used for VOC and COCO respectively. Specifically, we report $\text{AP}_{50}$ for VOC and $\{\text{AP},\text{AP}_{50},\text{AP}_{75},\text{AP}_{S},\text{AP}_{M},\text{AP}_{L}\}$ for COCO.
A.3 Hyperparameter Selections
@@
Appendix B More Ablation Studies
In this section, we conduct more ablation studies to further verify the effectiveness of the proposed DISCO. These ablation studies contain hyperparameter sensitivity, backbone compatibility, and the execution number of DISCO. Unless otherwise specified, the following experiments are all based on VOC at the $40\%$ noise level.
Hyper. Value $\text{AP}_{50}$bigstrut
$T$ 0.01 68.6
0.1 68.7
0.2 67.8
Hyper. Value $\text{AP}_{50}$bigstrut
[…]
10 68.7
[…]
Hyper. Value $\text{AP}_{50}$bigstrut
$\alpha$ 3 68.3
5 68.7
7 68.2
Hyper. Value $\text{AP}_{50}$bigstrut
$\beta$ 0.7 68.1
0.8 68.7
0.9 67.3
Hyper. Value $\text{AP}_{50}$bigstrut
[…]
0.3 68.7
[…]
Hyper. Value $\text{AP}_{50}$bigstrut
[…]
0.1 68.7
[…]
B.1 Hyperparameter Sensitivity
Here we evaluate the sensitivity of the hyperparameters used in DISCO. Note that we choose some moderate values rather than extreme ones to reasonably evaluate the sensitivity of each hyperparameter. The experimental results are reported in Table 5. As we can see, the temperature coefficient $T$ is relatively robust when set to 0.01 or 0.2. Tuning $T$ to a proper value can contribute to better performance. Besides, it can be observed that the augmented proposal number $N^{\prime}$ is insensitive when varying from $5$ to $20$. This is the reason why we empirically fix $N^{\prime}$ to $10$ for all settings. Moreover, the hyperparameters regulate the fusion of two bounding boxes (i.e., $\alpha$ and $\beta$) is also insensitive when varying within a moderate range, showing the effectiveness of our method. Furthermore, two loss weights $\gamma,\lambda$ also remain insensitive while $\lambda$ is relatively crucial. This is because it controls the strength of an extra classification loss term, directly affecting classification accuracy.
B.2 Backbone Compatibility
Following [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang], the benchmark experiments are performed with ResNet-50 [He et al.(2016)He, Zhang, Ren, and Sun] as the backbone. To further demonstrate the superior performance of our method, we conduct an additional experiment based on different backbones. Specifically, in this experiment, DISCO [Liu et al.(2022)Liu, Wang, Lu, Cao, and Zhang] is compared to OA-MIL on COCO at the $40\%$ noise level with the backbone set to ResNet-101 [He et al.(2016)He, Zhang, Ren, and Sun], and other experiment setups remain the same. In this way, we aim to evaluate the performance of our DISCO for a large-scale dataset when it is equipped with an advanced backbone. The experimental results are reported in Table 6. It can be observed that DISCO can further improve performance and still achieve SOTA results.
@@ B.2 Backbone Compatibility
Execution Num. $\text{AP}_{50}$bigstrut
@@
Table 6: Left: Ablation studies of backbone compatibility. The experiment is conducted on COCO at 40% noise level with ResNet-50 and ResNet-101. DISCO can still outperform OA-MIL when equipped with different backbones. Right: Ablation studies of the execution number of DISCO. Our execution strategy can achieve superior performance.
B.3 Execution Number of DISCO